# Quantitative Measures for Passive Sonar Texture Analysis

### v1

@@ 4.1 Synthetic and Real-World Datasets Evaluation
Table 3 summarizes the structural (StrTS) and statistical (StaTS) texture scores, along with classification accuracies for TDNN and HLTDNN across five datasets. The last column reports the accuracy improvement ($\Delta$ Accuracy) of HLTDNN over TDNN. The results support the hypothesis that HLTDNN, with its histogram-based representation, is better suited for statistically rich signals. In the statistical texture dataset, where StaTS is highest ($4.93\pm 0.04$), HLTDNN achieves a substantial +20.20% improvement over TDNN. This validates HLTDNN’s capacity to capture statistical variations more effectively than TDNN. The structural texture dataset exhibits the highest StrTS ($4.96\pm 0.03$) and a lower StaTS. While TDNN performs reasonably well ($72.50\pm 3.38$), HLTDNN still outperforms it with a +27.17% gain, demonstrating that its statistical modeling does not hinder its ability to capture structural patterns.
In the mixed statistical and structural textures dataset, where both texture types coexist, HLTDNN achieves near-perfect accuracy ($99.08\pm 0.88$) and shows a +26.90% gain over TDNN. This highlights HLTDNN’s adaptability in handling hybrid texture profiles. On the DeepShip dataset, which exhibits moderate statistical and lower structural texture complexity, HLTDNN outperforms TDNN by +8.06%. Despite being a real-world dataset with inherent variability, the gain shows HLTDNN’s effectiveness in generalizing to practical scenarios. VTUAD, the largest dataset, contains high values in both StrTS and StatS, and HLTDNN again shows a strong improvement of +10.53%. This result confirms the model’s scalability and ability to learn from complex, large-scale real-world data.
@@ 4.2 Pretraining Models Using Synthetic Datasets
Table 4: Model performance (%) on DeepShip with and without synthetic mixed texture pretraining. The best average accuracy is bolded.
Model Without Pretraining With Pretraining
TDNN $50.49\pm 0.74$ $52.5\pm 1.29$
HLTDNN $59.62\pm 1.69$ $62.11\pm 1.36$
To further investigate texture-aware learning, both models were pretrained on the mixed statistical and structural textures dataset and fine-tuned on DeepShip since the synthetic dataset was modeled after DeepShip’s characteristics. The performance comparison is shown in Table 4. Pretraining led to a consistent gain for both models, with HLTDNN improving by +2.49%. These results suggest that synthetic pretraining is particularly effective when the texture characteristics of the synthetic and real-world domains are well aligned. In summary, HLTDNN consistently outperforms TDNN across all datasets. The gains are especially notable in statistically textured conditions and further enhanced by texture-aligned pretraining. These findings highlight the benefit of explicitly modeling statistical texture and tailoring synthetic data to match real-world texture profiles.
@@ 5 Conclusion
This work investigated the role of structural and statistical texture in CNN-based passive sonar classification. Statistically rich signals frequently occur in passive sonar environments due to stochastic propagation, scattering, and background interference — phenomena that standard CNNs may fail to capture effectively. By constructing synthetic underwater datasets with controlled texture attributes, it was possible to isolate and analyze model behavior across varying degrees of signal complexity. Experimental findings show that traditional CNN architectures such as TDNN are more sensitive to structural regularities, whereas the texture-aware HLTDNN demonstrates improved performance on signals characterized by strong statistical texture.
These results support the importance of explicitly modeling statistical texture in passive acoustic target classification. The proposed statistical and structural texture quantification metrics further contribute to model interpretability by linking signal properties to performance characteristics. Future directions include generating synthetic data samples to adapt audio foundation models such as AudioPaLM [20] and AudioGPT [21] for passive sonar, analyzing texture sensitivity in transformer-based architectures such as the Audio Spectrogram Transformer (AST) [22], embedding texture-aware regularization into training objectives, leveraging more advanced approaches for data generation such as diffusion models [23] and applying the texture metrics to other audio domains.
6 Acknowledgment
Portions of this research were conducted with the advanced computing resources provided by Texas A&M High Performance Research Computing.

### v2

@@ 4.1 Synthetic and Real-World Datasets Evaluation
Table 3 summarizes the structural (StrTS) and statistical (StaTS) texture scores, along with classification accuracies for TDNN and HLTDNN across five datasets. The final column reports the absolute accuracy improvement ($\Delta$ Accuracy) of HLTDNN over TDNN. The results support the hypothesis that HLTDNN, with its histogram-based representation, is better suited for statistically rich signals. In the statistical texture dataset, where StaTS is highest ($4.93\pm 0.04$), HLTDNN achieves a substantial +20.20% improvement over TDNN, validating its ability to capture statistical variations more effectively. The structural texture dataset exhibits the highest StrTS ($4.96\pm 0.03$) and lower StaTS. While TDNN performs reasonably well ($72.50\pm 3.38$), HLTDNN still surpasses it with a +27.17% gain, indicating that statistical modeling does not hinder its ability to capture structural patterns.
In the mixed statistical and structural texture dataset, HLTDNN achieves near-perfect accuracy ($99.08\pm 0.88$) and shows a +26.90% gain over TDNN, demonstrating adaptability in handling hybrid texture profiles. On the DeepShip dataset, which presents moderate statistical and lower structural texture complexity, HLTDNN achieves an +8.06% improvement. Despite inherent variability in this real-world dataset, the performance gain reflects HLTDNN’s effectiveness in generalizing to practical scenarios. For the VTUAD dataset, which is the largest and exhibits high values in both StrTS and StaTS, HLTDNN delivers a strong +10.53% improvement. This confirms the model’s scalability and its ability to learn from complex, large-scale real-world data.
@@ 4.2 Pretraining Models Using Synthetic Datasets
Table 4: Performance (%) on DeepShip with pretraining using different synthetic datasets. Gains are reported relative to the baseline: TDNN = 50.49 ± 0.74, HLTDNN = 59.62 ± 1.69.
Pretraining Dataset TDNN HLTDNN
Accuracy Gain Accuracy Gain
Synthetic-Statistical 47.21 ± 1.89 –3.28 53.21 ± 1.89 –6.41
Synthetic-Structural 50.63 ± 1.95 +0.14 56.88 ± 1.22 –2.74
Synthetic Statistical and Structural 52.50 ± 1.29 +2.01 62.11 ± 1.36 +2.49
To evaluate the impact of synthetic pretraining, both models were pretrained on synthetic datasets and then fine-tuned on DeepShip. As shown in Table 4, pretraining with the mixed texture dataset improved performance for both TDNN and HLTDNN, with a gain of +2.01% and +2.49%, respectively. In contrast, statistical-only or structural-only pretraining resulted in minimal or negative gains. This indicates that synthetic pretraining is particularly effective when the texture characteristics of the synthetic and real-world domains are well aligned. The mixed synthetic statistical & structural dataset, which captures both statistical and structural properties, more closely reflects the texture complexity present in DeepShip, leading to better generalization after pretraining.
@@ 5 Conclusion
This work explored the role of structural and statistical texture in CNN-based passive sonar classification. Statistically rich signals, common in passive sonar due to stochastic propagation and interference, are often underrepresented in standard CNN processing. Synthetic datasets with controlled texture characteristics enabled systematic evaluation of model performance under varying texture complexities. Results show that while TDNN favors structural regularities, the texture-aware HLTDNN performs better on statistically textured signals, highlighting the value of modeling statistical texture. The proposed texture quantification metrics offer interpretability by linking signal properties to model behavior. Future work includes adapting foundation models like AudioPaLM [20] and AudioGPT [8], evaluating transformer-based models such as AST [6], integrating texture-aware regularization, using diffusion-based generation [11], and extending this framework to other audio domains.